OpenStudy (anonymous):

-2x2 - 16x - 44 = 0. x equals negative 8 plus or minus I square root of 6 x equals negative 8 plus or minus 2i square root of 6 x equals negative 4 plus or minus i square root of 6 x equals negative 4 plus or minus 2i square root of 6

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

The Quadratic formula\[ax^2+bx+c=0\]\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

it b right?

OpenStudy (unklerhaukus):

i dont know